Find the modulus and the principal argument of each of the given complex numbers.
(a) 3 − 4i, (b) −2 + i, (c) 1
1 + i

3
, (d) 7 − i
−4 − 3i
,
(e) 5(cos π/3 + isin π/3), (f) cos 2π/3 − sin 2π/3.
1
Expert's answer
2020-06-03T18:57:16-0400

We will use the notation "z = a+ib" for given complex number. So, modulus of "z" is

"|z| = \sqrt{a^2+b^2}" and principal argument of "z" is "arg(z) = tan^{-1}(\frac{b}{a}) \in (-\pi,\pi)."

a) Given "z = 3-4i"

Hence, "a = 3, b = -4".

"\implies |z| = \sqrt{3^2+(-4)^2} = 5"

And, since z lies in fourth quadrant so "arg(z) = - tan^{-1} (\frac{4}{3})".


b) Given "z = -2+i" .

So, "|z| = \sqrt{2^2+1^2} = \sqrt{5}"

and z lies in 2nd quadrants so "arg(z) = \pi - tan^{-1}(\frac{1}{2})".


c) Given "z = 1+i\sqrt{3}" .

Hence, "|z| = \sqrt{1^2+3} = \sqrt{4} = 2"

and z lies in first quadrant so "arg(z) = tan^{-1}(\sqrt{3}) = \frac{\pi}{3}"


d) Given "z = -4-3i".

Hence, "|z| = \sqrt{4^2+3^2} = \sqrt{25} = 5"

and z lies in 3rd quadrants, so "arg(z) = \pi+tan^{-1}(\frac{3}{4})".


e) Given "z = 5(cos (\u03c0\/3) + isin( \u03c0\/3))"

Hence, "|z| = \sqrt{5^2(cos^2(\pi/3)+sin^2(\pi/3))} = 5"

And "arg(z) = tan^{-1}(\frac{5sin(\pi/3)}{5cos(\pi/3)}) = tan^{-1} (tan(\pi/3)) = \pi/3" .


f) Given "z = cos(2\pi/3) - sin(2\pi/3) = -cos(\pi/3)-sin(\pi/3) = - \frac{1}{2} - \frac{\sqrt{3}}{2}".

So, "|z| = \sqrt{( cos(2\pi/3) - sin(2\pi/3))^2} = cos(2\pi/3) - sin(2\pi/3) = - (\frac{1+\sqrt{3}}{2})"

and given z lies on negative "x-"axis, hence "arg(z) = tan^{-1}(0) = -\pi" .